Smithfield and Whoville are 184 miles apart. A car leaves Smithfield travelling towards Whoville at the same time another car leaves Whoville travelling towards Smithfield. The Car leaving Smithfield averages 10 mph more than the other and they meet after 1 hour and 36 minutes. What are the average speeds of the cars?

Answer by checkley77(12844)

D=RT
184=(X+X+10)1.6
184=(2X+10)1.6
184=3.2X+16
3.2X=184-16
3.2X=168
X=168/3.2
X=52.5 MPH. FOR THE SLOWER CAR.
52.5+10=62.5 MPH. FOR THE FASTER CAR.
PROOF:
`84=(52.5+62.5)1.6
184=115*1.6
184=184